# Electric-Field Driven Nuclear Dynamics of Liquids and Solids from a Multi-Valued Machine-Learned Dipolar Model

### v1

@@ III.1 Dielectric Properties and Electrofreezing of Water
For the classical-nuclei simulations, we also calculated the water self-diffusion constant at different $\mathcal{E}$, shown in Fig. 2d. Consistent with the phenomenon of electrofreezing [23], we observe this coefficient to decrease from around 1.2 Å2/ps at zero-field to 0.3 Å2/ps at $\mathcal{E}=0.15$ V/Å. Based on the previous considerations, it is certain that NQE will also enhance this effect by causing a faster decrease in the diffusion coefficient with increasing field strength.
III.2 Phase Transition and Non-equilibrium Phonon Driving in LiNbO3
@@
We have evaluated the mean and the variance (fluctuations) of the dipole in the temperature range of $10$ to $1800$ K. The results are shown in Fig. 3a and b. Both plots show the typical behavior of a second-order phase transition [26]. The average value of the order parameter (the dipole) decreases to zero above the Curie temperature (Fig. 3a), while its fluctuations show a sharp peak at the same temperature (Fig. 3b). Finite size effects cause this peak to broaden with respect to the expected $\delta$-function. The analysis of the atomic displacements presented in Supplementary Section S4, which are in agreement with simulations reported in Ref. [59], confirm that the simulations indeed capture the ferroelectric to paraelectric phase-transition.
@@
The phase-transition temperature inferred from Fig. 3a and b matches quite well the experimental Curie temperature $T_{C}$, which is likely to be partially fortuitous, because our simulations do not include the effects of thermal expansion. These effects are expected to increase $T_{C}$ by around 100 K [59]. Nevertheless, the reasons why the value we calculate is so close to the experimental result are that we properly account for all orders of anharmonic coupling between vibrational modes and that we accurately model its polarization changes, including its multi-valued nature. The absence of anharmonic couplings was previously shown to cause an underestimation of 250-300 K for $T_{C}$ [59, 60].
@@
LiNbO3 has been at the center of active scientific debates due to the possibility of reversing its polarization using THz laser pulses which induce a non-thermal excitation of an IR-active mode $Q_{\rm IR}$ that strongly couples to a polarization-reversal mode $Q_{\rm P}$, as theoretically proposed in Ref. [61] and partially realized in Ref. [28]. To date, theoretical studies modeling this effect [62, 63, 29] were based on approximations to the PES which consider two or only a few phonon modes and the lowest-order anharmonic couplings between them. We have employed the ML models developed in this work to address this question, without relying on such approximations.
@@
We have performed simulations where we applied a time-dependent monochromatic pulse with frequency $\nu=18$ THz enveloped with a Gaussian function with FWHM=188 fs and different maximum field intensities $\mathcal{E}_{\mathrm{max}}$ in the range $0.05-0.30$ V/Å. These values were chosen based on the reported experimental setups [28, 29] (see also Supplementary Section S5). The excitation is in close resonance with the 18.4 THz $Q_{\rm IR}$ mode, while the proposed $Q_{\rm P}$ mode lies at $7.4$ THz. These modes are depicted in Supplementary Section S6.
@@
As shown in the Supplementary Sections S6 and S7, we observe the non-linear coupling of the $Q_{\rm IR}$ mode that is directly excited by the pulse at $\nu=18$ THz with the $Q_{\rm P}$ mode, which starts accumulating energy with a delay of some hundreds of fs. In addition we observe that only two other modes $\Gamma$-point modes, of A1 symmetry, at 8.2 and 10.0 THz get excited at later times. These modes are also IR active and, in particular the mode at 8.2 THz contributes to the polarization-reversal, attesting to the complex dynamical coupling in this system. In Fig. 3c we show the time-dependent variation of the polarization at times previous to, during (gray shaded area) and after the laser pulse is applied. The value of the polarization deviates from the thermal equilibrium value as soon the laser pulse starts acting on the system. We can recognize three different regimes, depending on the maximum intensity of the laser pulse. (i) At low intensities such as $0.05$ V/Å, the polarization is only slightly perturbed and the system remains ferroelectric. (ii) At higher intensities ($0.15$ and $0.20$ V/Å), the system is transiently driven to the paraelectric state (with zero polarization) during the laser pulse, but returns to the original ferroelectric state with a relaxation time of about $300$ fs, during which coherent dynamics can still be observed. (iii) At extremely high fields ($0.25$ and $0.30$ V/Å), the system transiently reaches higher values of reversed polarization but never fully switches, and after the pulse is switched off it transitions to the paraelectric state with incoherent dynamics because the amount of energy absorbed effectively heats the system above $T_{C}$.
@@ IV Conclusion
V Methods
The evolution of the nuclear equations of motion including time-dependent and independent applied electric fields was obtained by an implementation developed by us in the i-PI code, which was recently described in Ref. [66]. We refer the reader to that publication for technical details. All molecular-dynamics simulations were run with i-PI as the driving code communicating with the ML models, which provided all components of the force, and which are further described below.
@@ V Methods
Table 1: Unit cell parameters used for the calculation of the oxidation numbers as shown in Section II.1.
@@ H $\left[0,\phantom{-}1,\phantom{-}0\right]$ 2 -1
Table 2: Computational details used to evaluate the oxidation numbers (as presented in Section II.1) of all the atomic species in water and LiNbO3. The table reports the direction, on the basis of the lattice vectors, along which the atoms were displaced, the total number of atoms of that species, and the evaluated oxidation number.
@@
V.1 Water
@@
We have evaluated the oxidation numbers of the H and O atoms of a single water molecule following Section II.1. A cubic simulation box of $4$  (whose parameters are reported in Table 1) was used. Figure 1b was obtained considering $100$ configurations, where the H atoms were uniformly displaced along the direction reported in Table 2 from the relaxed initial configuration, while the O atom was fixed, and vice-versa. The DFT calculations were performed with the revPBE functional, no spin polarization, a $3\!\times\!3\!\times\!$ 3 × 3 $\bm{k}$-grid, atomic ZORA relativistic treatment [67], and the intermediate basis set in the FHI-aims [68] code. The polarization components were calculated [69] on a denser $\bm{k}$-grid, namely $40\!\times\!10\!\times\!10$ for the component along $\bm{a}_{1}$, $10\!\times\!40\!\times\!10$ for $\bm{a}_{2}$, and $10\!\times\!10\!\times\!40$ for $\bm{a}_{3}$.
@@
The machine-learning interatomic potential (MLIP) for liquid water was trained with a MACE [44, 45] model on the data provided in Ref [70], i.e. energy and forces computed using the revPBE0 functional with Grimme D3 [71, 72] dispersion corrections. We have trained a MACE MLIP with a cutoff radius of $6$, $2$ layers, $64$ embedding channels, $l=3$ as the highest spherical harmonics, a maximum of $l=0$ for each message, and correlation order $3$ at each interaction layer.
@@
The dataset for training the dipole model was obtained from 3 NVT simulations (stochastic-velocity-rescaling thermostat SVR  [73]; $\tau=100$ fs; $T=330,360,390$ K) with i-PI [66] and the trained MLIP as a force provider. We ran $100$ ps for each run with a time-step of $1$ fs. The simulation box was cubic with lattice vectors of length $14.9365$  and $111$ molecules, corresponding to a density of $0.9965$ g/m-3. The farthest point sampling (FPS) algorithm, as implemented in the librascal library [74] and based on SOAP descriptors as reported in Ref. [75] was used to select the $1000$ most diverse structures from all trajectories.
@@
We used the structures in the unit-cell in an aperiodic setting (effectively water droplets) and computed the dipole using Eq. (4) with the revPBE functional. This was done solely to increase efficiency of data generation. We have split the data with a ratio of $80/20$ for training and validation, after having separated 20% of the full dataset for testing. The dipole model was trained using a modified MACE architecture developed by us and available in Ref. [76], with a cutoff radius of $6$, $2$ layers, $14\!\times\!0e\!+\!4\!\times\!$ 1o as irreps for the hidden node states, $l=2$ as the highest spherical harmonic, correlation order of $3$ at each layer. We added the oxidation numbers contribution according to Eq. (9) in the training. The performance of these models are reported in Table 3.
@@
The IR absorption spectra of liquid water and field-dependent self-diffusion coefficients (Fig. 2a and d) were obtained by first equilibrating the system for $100$ ps at $300$ K (Langevin thermostat, $\tau=50$ fs, $dt=0.25$ fs) and then applying the constant electric fields following Eqs. (1) and 2. The MLIP together with the ML dipole model were used to obtain the forces. We picked $16$ random snapshots from this trajectory and used them to start 16 NVE trajectories of $40$ ps each. The IR absorption spectra were obtained from the Fourier transform of the autocorrelation of the dipole derivatives [77]. The diffusion coefficient was obtained by computing the mean square displacement (MSD) of the O atoms for each time-step and taking the average value of its time-derivative on the last $10$ ps, when the behavior is already linear with time.
@@
The high-frequency part of the dielectric susceptibility was computed with the fluctuation-dissipation relation of the dipole and dipole-derivative equilibrium correlation function [51, 52, 10], using the same data as for the infrared absorption spectrum with zero electric field. The low-frequency part of the dielectric susceptibility was computed using the non-equilibrium direct electric field method described in Ref. [51]. We simulated NVT trajectories of $1-2$ ns (SVR thermostat, $\tau=100$ fs, $dt=0.5$ fs) including a time-dependent monochromatic external electric field of $\mathcal{E}_{\rm max}=0.04$ V/Å, with frequencies of $1,3,10,30,100$ and $200$ GHz. The time-dependent behavior of the (induced) polarization at each frequency was fitted with a function $P_{\rm ind}\cos\left(2\pi\nu-\phi\right)$ to obtain the parameters $P_{\rm ind}$ and $\phi$. The introduction of the 3rd and 5th harmonics in the fit, as done in Ref. [51], resulted in negligible corrections to the fitted values. The real and imaginary parts of the dielectric susceptibility were then computed as $P_{\rm ind}\cos\left(\phi\right)/\varepsilon_{0}\mathcal{E}_{\rm max}$ and $P_{\rm ind}\sin\left(\phi\right)/\varepsilon_{0}\mathcal{E}_{\rm max}$ respectively.
@@
The dipole and structural data in Figs. 2c and d were derived from $16$ NVT trajectories for each different electric field of $10$ ps each. The static structural properties related to electrofreezing were also calculated including nuclear quantum effects with path integral molecular dynamics (PIMD). We ran the same number and length of trajectories for each electric field intensity, for 30 ps, using $8$ ring-polymer replicas, a time-step of 0.25 fs, and the PIGLET thermostat [78].
V.2 LiNbO3

### v2

@@ II.2 Dielectric Properties and Electrofreezing of Water
We also calculated the water self-diffusion coefficient $D_{\mathrm{self}}$ at different $\mathcal{E}$, shown in Fig. 2d. Consistent with the phenomenon of electrofreezing [26], we observe this coefficient to decrease from around 0.20 Å2/ps at zero-field to 0.05 Å2/ps at $\mathcal{E}=0.15$ V/Å in the classical case. Consistent with the previous considerations, NQE also enhance this effect, causing a faster decrease in $D_{\mathrm{self}}$ with increasing field strength. At zero-field, previous reports have shown that NQE increase $D_{\mathrm{self}}$ with respect to the classical simulations by about 12% [54], and here we show that it is instead smaller than the classical value at $\mathcal{E}=0.15$ V/Å, as shown in Fig. 2d.
II.3 Phase Transition and Non-equilibrium Phonon Driving in LiNbO3
@@
We have evaluated the mean and the variance (fluctuations) of the dipole in the temperature range of $10$ to $1800$ K. The results are shown in Fig. 3a and Fig. 3b. Both plots show the typical behavior of a second-order phase transition [30]. The average value of the order parameter (the dipole) decreases to zero above the Curie temperature (Fig. 3a), while its fluctuations show a sharp peak at the same temperature (Fig. 3b). Finite size effects cause this peak to broaden with respect to the expected $\delta$-function. The analysis of the atomic displacements presented in Supplementary Section S5, which are in agreement with simulations reported in Ref. [56], confirm that the simulations indeed capture the ferroelectric to paraelectric phase-transition.
@@
The phase-transition temperature inferred from Fig. 3a and Fig. 3b matches quite well the experimental Curie temperature $T_{C}$, which is likely to be partially fortuitous, because our simulations do not include the effects of thermal expansion. These effects are expected to increase $T_{C}$ by around 100 K [56]. Nevertheless, the reasons why the value we calculate is so close to the experimental result are that we properly account for all orders of anharmonic coupling between vibrational modes and that we accurately model its polarization changes, including its multi-valued nature. The absence of anharmonic couplings was previously shown to cause an underestimation of 250-300 K for $T_{C}$ [56, 57].
@@
LiNbO3 has been at the center of active scientific debates due to the possibility of reversing its polarization using THz laser pulses which induce a non-thermal excitation of an IR-active mode $Q_{\rm IR}$ that strongly couples to a polarization-reversal mode $Q_{\rm P}$, as theoretically proposed in Ref. [58] and partially realized in Ref. [32]. To date, theoretical studies modeling this effect [59, 60, 33] were based on approximations to the potential energy surface (PES) which consider two or only a few phonon modes and the lowest-order anharmonic couplings between them. We have employed the ML models developed in this work to address this question, without relying on such approximations.
@@
We have performed simulations where we applied a time-dependent monochromatic pulse with frequency $\nu=18$ THz enveloped with a Gaussian function with FWHM=188 fs and different maximum field intensities $\mathcal{E}_{\mathrm{max}}$ in the range 0.050.300.05-0.30 V/Å. These values were chosen based on the reported experimental setups [32, 33] (see also Supplementary Section S6). The excitation is in close resonance with the 18.4 THz $Q_{\rm IR}$ mode, while the proposed $Q_{\rm P}$ mode lies at $7.4$ THz. These modes are depicted in Supplementary Section S7.
@@
As shown in the Supplementary Sections S7 and S8, we observe the non-linear coupling of the $Q_{\rm IR}$ mode that is directly excited by the pulse at $\nu=18$ THz with the $Q_{\rm P}$ mode, which starts accumulating energy with a delay of some hundreds of fs. In addition we observe that only two other $\Gamma$-point modes, of A1 symmetry, at 8.2 and 10.0 THz get excited at later times. These modes are also IR active and, in particular the mode at 8.2 THz contributes to the polarization-reversal, attesting to the complex dynamical coupling in this system. In Fig. 3c we show the time-dependent variation of the polarization at times previous to, during (gray shaded area) and after the laser pulse is applied. The value of the polarization deviates from the thermal equilibrium value as soon the laser pulse starts acting on the system. We can recognize three different regimes, depending on the maximum intensity of the laser pulse. (i) At low intensities such as $0.05$ V/Å, the polarization is only slightly perturbed and the system remains ferroelectric. (ii) At higher intensities ($0.15$ and $0.20$ V/Å), the system is transiently driven to the paraelectric state (with zero polarization) during the laser pulse, but returns to the original ferroelectric state with a relaxation time of about $300$ fs, during which coherent dynamics can still be observed. (iii) At extremely high fields ($0.25$ and $0.30$ V/Å), the system transiently reaches higher values of reversed polarization but never fully switches, and after the pulse is switched off it transitions to the paraelectric state with incoherent dynamics because the amount of energy absorbed effectively heats the system above $T_{C}$.
@@ III Discussion
IV Methods
IV.1 Theory of Molecular Dynamics Driven by Electric Fields
[…]
where $T(\bm{p})$ is the kinetic energy, $\bm{p}$ the nuclear momenta, $\bm{R}$ the nuclear coordinates, $V(\bm{R})\equiv E[n_{0}]$ is the Born-Oppenheimer (BO) PES obtained from the ground-state electronic density $n_{0}$, and $\bm{\mu}(\bm{R})\equiv\bm{\mu}[n_{0}]$ is the system dipole.
[…]
The first term on the right-hand side is the usual BO force. The second term can be expressed in terms of the atomic-polar tensors [20, 63], also called BEC $\bm{Z}^{*}$,
[…]
For a given atomic configuration, the BEC can be computed from any of the two expressions in Eq. (8). The (relative) efficiency of these approaches strongly depends on the system, on the chosen electronic-structure code, and on details of the implementations. Evidently, the expressions can be calculated by finite differences or by perturbation theory [64, 65]. While the latter is elegant and can be quite efficient especially for systems where symmetry can be exploited, the former ensures a natural handling of electronic-structure level of approximation, such as different density-functionals, relativistic corrections and other terms that are cumbersome to evaluate analytically. In any event, the calculation of $\bm{Z}^{*}$ always represents an added cost on the already costly electronic-structure calculation of the dipole.
We circumvent the expensive first-principles calculation of $\bm{Z}^{*}$ that enter Eq. (7) by developing a ML model that is trained only on $\bm{\mu}$ and that is autodifferentiable [66] with respect to its input quantities, the nuclear coordinates $\bm{R}$. Training on $\bm{\mu}$ presents some advantages, including the ease to produce data for this quantity with diverse electronic-structure methods and the guarantee that the acoustic sum rule [44] for $\bm{Z}^{*}$ in the model is preserved, making it strictly translationally invariant.
The evolution of the nuclear equations of motion including time-dependent and independent applied electric fields was obtained by an implementation developed by us in the i-PI code, which was recently described in Ref. [67]. We refer the reader to that publication for technical details. All molecular-dynamics simulations were run with i-PI as the driving code communicating with the ML models, which provided all components of the force.
@@ IV.1 Theory of Molecular Dynamics Driven by Electric Fields
Table 1: Unit cell parameters used for the calculation of the oxidation numbers as shown in Section IV.2.
@@ H $\left[0,\phantom{-}1,\phantom{-}0\right]$ 2 -1
Table 2: Computational details used to evaluate the oxidation numbers (as presented in Section IV.2) of all the atomic species in water and LiNbO3. The table reports the direction, on the basis of the lattice vectors, along which the atoms were displaced, the total number of atoms of that species, and the evaluated oxidation numbers.
@@
IV.2 Dipoles in Aperiodic and Periodic Systems
$\bm{\mu}_{e}=-e\int_{\mathbb{R}^{3}}d\bm{r}\,n_{0}(\bm{r})\bm{r}.$ (9)
In periodic, condensed-phase systems, $\bm{\mu}_{e}$ cannot be directly evaluated by the expression above, as restricting the integration domain to the primitive unit cell would yield boundary-sensitive results. The dipole is related to the polarization $\bm{P}$ of a periodic system, in that the polarization is the dipole per unit volume, $\bm{P}=\bm{\mu}/\Omega$. The modern theory of polarization (MTP) [68, 69, 70] provides a proper definition of this quantity in periodic systems, showing that the electronic contributions to $\bm{P}$ are related to a geometric phase of the wavefunction. There are several different (but equivalent) ways to evaluate $\bm{\mu}_{e}$ within MTP, which are all related to a way of evaluating the Berry phase [69, 70, 71]. As a consequence, the components of $\bm{\mu}$ corresponding to each lattice vector $\bm{a}_{\alpha}$ can be defined only up to quantum of polarization $e\left|\bm{a}_{\alpha}\right|$ and the dipole along $\bm{a}_{\alpha}$ is
$\mu_{\alpha}\equiv\mu_{\alpha}\mod e|\bm{a}_{\alpha}|.$ (10)
IV.3 Water Simulation Details
@@
We have evaluated the oxidation numbers of the H and O atoms of a single water molecule following Section IV.2. A cubic simulation box of $4$ Å  (whose parameters are reported in Table 1) was used. Figure 1b was obtained considering $100$ configurations, where the H atoms were uniformly displaced along the direction reported in Table 2 from the relaxed initial configuration, while the O atom was fixed, and vice-versa. The DFT calculations were performed with the revPBE functional, no spin polarization, a $3\!\times\!3\!\times\!3$ $\bm{k}$-grid, atomic ZORA relativistic treatment [72], and the intermediate basis set in the FHI-aims [73] code. The polarization components were calculated [74] on a denser $\bm{k}$-grid, namely $40\!\times\!10\!\times\!10$ for the component along $\bm{a}_{1}$, $10\!\times\!40\!\times\!10$ for $\bm{a}_{2}$, and $10\!\times\!10\!\times\!40$ for $\bm{a}_{3}$.
@@
The MLIP for liquid water was trained with a MACE [39, 40] model on the data provided in Ref [75], i.e. energy and forces computed using the revPBE0 functional with Grimme D3 [76, 77] dispersion corrections. We have trained a MACE MLIP with a cutoff radius of $6$ Å, $2$ layers, $64$ embedding channels, $\ell=3$ as the highest spherical harmonics, a maximum of $L=0$ for each message, and correlation order $3$ at each interaction layer.
@@
The dataset for training the dipole model was obtained from 3 NVT simulations (stochastic-velocity-rescaling thermostat SVR  [78]; $\tau=100$ fs; $T=330,360,390$ K) with i-PI [67] and the trained MLIP as a force provider. We ran $100$ ps for each run with a time-step of $1$ fs. The simulation box was cubic with lattice vectors of length $14.9365$ Å  and $111$ molecules, corresponding to a density of $0.9965$ g/m-3. The farthest point sampling (FPS) algorithm, as implemented in the librascal library [79] and based on SOAP descriptors as reported in Ref. [80] was used to select the $1000$ most diverse structures from all trajectories.
@@
We used the structures in the unit-cell in an aperiodic setting (effectively water droplets) and computed the dipole using Eq. (9) with the revPBE functional. This was done solely to increase efficiency of data generation. We have split the data with a ratio of $80/20$ for training and validation, after having separated 20% of the full dataset for testing. The dipole model was trained using a modified MACE architecture developed by us and available in Ref. [81], with a cutoff radius of $6$ Å, $2$ layers, $4\!\times\!0e\!+\!4\!\times\!1$o as irreps for the hidden node states, $\ell=2$ as the highest spherical harmonic, correlation order of $3$ at each layer. We added the oxidation numbers contribution according to Eq. (4) in the training. The performance of these models are reported in Table 3.
@@
The IR absorption spectra of liquid water and field-dependent self-diffusion coefficients (Fig. 2a and d) were obtained by first equilibrating the system for $100$ ps at $300$ K (Langevin thermostat, $\tau=50$ fs, $dt=0.25$ fs) and then applying the constant electric fields following Eqs. (6) and 7. The MLIP together with the ML dipole model were used to obtain the forces. We picked $16$ random snapshots from this trajectory and used them to start 16 NVE trajectories of $40$ ps each. The IR absorption spectra were obtained from the Fourier transform of the autocorrelation of the dipole derivatives [82]. The diffusion coefficient was obtained by computing the mean square displacement (MSD) of the O atoms for each time-step and taking the average value of its time-derivative on the last $10$ ps, when the behavior is already linear with time
@@
The high-frequency part of the dielectric susceptibility was computed with the fluctuation-dissipation relation of the dipole and dipole-derivative equilibrium correlation function [47, 48, 13], using the same data as for the infrared absorption spectrum with zero electric field. The low-frequency part of the dielectric susceptibility was computed using the non-equilibrium direct electric field method described in Ref. [47]. We simulated NVT trajectories of 1 ns (SVR thermostat, $\tau=100$ fs, $dt=0.5$ fs) including a time-dependent monochromatic external electric field of $\mathcal{E}_{\rm max}=0.04$ V/Å, with frequencies of $1,3,10,30,100$ and $200$ GHz. The time-dependent behavior of the (induced) polarization at each frequency was fitted with a function $P_{\rm ind}\cos\left(2\pi\nu-\phi\right)$ to obtain the parameters $P_{\rm ind}$ and $\phi$. The introduction of the 3rd and 5th harmonics in the fit, as done in Ref. [47], resulted in negligible corrections to the fitted values. The real and imaginary parts of the dielectric susceptibility were then computed as $P_{\rm ind}\cos\left(\phi\right)/\varepsilon_{0}\mathcal{E}_{\rm max}$ and $P_{\rm ind}\sin\left(\phi\right)/\varepsilon_{0}\mathcal{E}_{\rm max}$ respectively.
@@
The dipole and structural data in Figs. 2c and d were derived from $16$ NVT trajectories for each different electric field of $10$ ps each. The static structural properties related to electrofreezing were also calculated including nuclear quantum effects with PIMD. We ran the same number and length of trajectories for each electric field intensity, for 30 ps, using $8$ ring-polymer replicas, a time-step of 0.25 fs, and the PIGLET thermostat [83].
IV.4 LiNbO3 Simulation Details